# Automated Data Augmentation for Few-Shot Time Series Forecasting: A Reinforcement Learning Approach Guided by a Model Zoo

### v2

@@ 5.2 Baseline Augmentation Methods
We employ the Anchor Data Augmentation (ADA) method, which improves the domain-agnostic Mixup techniques. ADA uses multiple replicas of modified samples generated by Anchor Regression (AR) to create additional training examples.
@@ 5.3 Results of Few-Shot Time Series Forecasting
Under the few-shot learning setup, we evaluate the effectiveness of ReAugment on various forecasting models, including iTransformer (Liu et al., 2024b), PatchTST (Nie et al., 2023), and DLinear (Zeng et al., 2023). As illustrated in Table 1, on most public datasets, applying our method to construct a model zoo and perform automated data augmentation consistently improves the prediction accuracy of these two popular time series forecasting approaches. This demonstrates that our method is adaptable to different forecasting models and can effectively enhance data augmentation.
In Table 2, we compare the performance of different augmentation methods. All augmentation methods use iTransformer as the forecasting model and expand the original training set by three times. Notably, ReAugment consistently outperforms all methods across all datasets, including the state-of-the-art ADA method. In most cases, however, data-agnostic augmentation methods, such as Gaussian and Convolve, result in a negative impact on performance.
@@ 5.5 Model Analyses
Evaluating the improved forecasting accuracy achieved through data augmentation provides a direct measure of our method’s effectiveness. However, it does not account for the fact that the impact of data scarcity can vary significantly across different datasets. We propose a new metric that calculates the ratio of the performance promotion achieved through data augmentation in the few-shot setup, compared to the improvement obtained using the full training set. For example, when using MSE, it can be formulated as:
This metric considers the nature of the dataset and provides a more comprehensive evaluation of the value of augmentation. As shown in Table 6, a larger value indicates a greater performance improvement due to data augmentation.
| Metric | ETTh1 | ETTh2 | ETTm1 | ETTm2 | Average |
| 41.4% | 191.7% | 47.6% | 70.0% | 87.7% | |
| 33.3% | 94.7% | 26.4% | 44.4% | 49.7% | |
| Weather | Elec | Traffic | Exc | ||
| 16.7% | 21.1% | 51.0% | 18.2% | 26.8% | |
| 22.2% | 15.0% | 50.0% | 35.3% | 30.6% |
@@
Inspired by preliminary findings, we augment the top overfit-prone samples. What would be the impact of augmenting more or fewer samples? To investigate this, we compare augmenting the top and top overfit-prone samples with augmenting the entire few-shot dataset. For consistency, the total amount of augmented data was kept three times the size of the original training set and applied to the iTransformer model. As shown in Table 5, directly augmenting the entire few-shot dataset was less effective than data-dependent augmentation, which aligns with our preliminary findings. Furthermore, augmenting different percentages of high-variance samples resulted in varying degrees of improvement, demonstrating that our data augmentation model can adaptively enhance overfit-prone samples, leading to better performance of the forecasting model.
@@ 6 Conclusions and Limitations
In this paper, we proposed ReAugment, a novel data augmentation method driven by reinforcement learning. There are two key technical contributions in methodology: First, our method automatically identifies the critical training data for augmentation, termed the overfit-prone samples, based on prediction diversity from a set of pretrained models. Second, ReAugment exploits a variational masked autoencoder in conjunction with the REINFORCE algorithm to generate new data from these overfit-prone samples. ReAugment significantly boosts forecasting performance while maintaining minimal computational overhead by leveraging a learnable policy to transform the overfit-prone samples.
One unresolved issue in this study is the reliance on multiple pretrained models, which can increase the computational complexity of the approach. While we briefly discuss the computational cost in the appendix, this limitation is important to consider for practical applications.
Impact Statement
The potential social impact of our data augmentation method for few-shot time series forecasting could be substantial, especially in areas like economics, healthcare, climate prediction, and other fields that rely on time-sensitive forecasting. Our method could dramatically improve the accuracy of forecasts in situations where data is scarce. In industries like finance or energy, where historical data might be limited or costly to gather, being able to produce more reliable predictions with minimal data would have a significant impact. This could reduce risk and uncertainty, leading to more informed decisions.
While our method could have many benefits, it also raises some ethical questions. First, if the augmentation process isn’t well-calibrated, it could introduce biases that lead to inaccurate or unfair predictions. For instance, if certain data patterns are overly amplified, predictions in specific regions or for certain demographics could be skewed. Second, in sensitive domains like healthcare, time series data augmentation could be used to generate predictions without compromising patient privacy. However, there must be a balance between improving predictive accuracy and safeguarding individual privacy.
Appendix
@@ Appendix A Overall Training Pipeline
We present the pseudocode of the overall training pipeline in Algorithm 1.
@@ Appendix B Dataset Details
ETT consists of two hourly-level datasets (ETTh) and two -minute-level datasets (ETTm). Each of them contains factors of electricity transformers including load and oil temperature from July 2016 to July 2018.
@@
For the standard setup, we follow the data processing method of iTransformer (Liu et al., 2024b), dividing the dataset into training, validation, and test sets, with this partitioning strictly aligned in chronological order.
@@
Appendix C Implementation Details
C.1 Hyperparameters
@@
C.2 Computing Resources and Computational Costs
We perform all experiments on an NVIDIA RTX 3090 GPU. The proposed data augmentation method involves training VMAE and employing the REINFORCE algorithm, and it requires backtesting across multiple models within the model zoo. These processes introduce additional computational overhead compared to basic prediction models. As shown in Table 9, we report the training time for VMAE (Stage A), REINFORCE (Stage B), and the forecasting model (Stage C) under the few-shot setup. The increased training time (including Stage A and Stage B) is considered acceptable due to the significant performance gains. Besides, it is notably shorter than the time needed for training the forecasting models.
| Dataset | Stage A: VMAE | Stage B: REINFORCE | Stage C: Forecasting | |
| iTransformer | PatchTST | |||
| ETTh1 | 1min | 2min | 2min | 5min |
| Elec. | 24min | 31min | 1h 33min | 2h 24min |
| Traffic | 1h 22min | 1h 53min | 4h 27min | 6h 50min |

### v3

@@ 5.1 Experimental Setups
ADA [26]: We employ the Anchor Data Augmentation (ADA) method, which improves the domain-agnostic Mixup techniques. ADA uses multiple replicas of modified samples generated by Anchor Regression to create additional training examples.
| Dataset | Original | Gaussian | Convolve | TimeGAN | ADA | ReAugment | ||||||
| MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | |
| ETTh1 | 0.434 | 0.411 | 0.437 | 0.416 | 0.441 | 0.417 | 0.444 | 0.419 | 0.435 | 0.413 | 0.4220.01 | 0.4030.01 |
| ETTh2 | 0.362 | 0.320 | 0.365 | 0.321 | 0.364 | 0.323 | 0.366 | 0.327 | 0.368 | 0.331 | 0.3390.01 | 0.3020.01 |
| ETTm1 | 0.440 | 0.470 | 0.438 | 0.469 | 0.426 | 0.479 | 0.430 | 0.483 | 0.429 | 0.484 | 0.4100.01 | 0.4360.01 |
| ETTm2 | 0.282 | 0.204 | 0.283 | 0.204 | 0.286 | 0.207 | 0.285 | 0.206 | 0.284 | 0.207 | 0.2750.02 | 0.1960.01 |
| Weather | 0.231 | 0.187 | 0.240 | 0.196 | 0.253 | 0.204 | 0.239 | 0.191 | 0.246 | 0.198 | 0.2290.00 | 0.1850.00 |
| Electricity | 0.258 | 0.168 | 0.263 | 0.170 | 0.262 | 0.170 | 0.267 | 0.177 | 0.265 | 0.171 | 0.2540.01 | 0.1650.01 |
| Traffic | 0.318 | 0.466 | 0.319 | 0.467 | 0.320 | 0.463 | 0.315 | 0.449 | 0.318 | 0.456 | 0.2930.01 | 0.4290.01 |
| Exchange | 0.228 | 0.103 | 0.229 | 0.104 | 0.226 | 0.099 | 0.226 | 0.100 | 0.235 | 0.116 | 0.2240.00 | 0.0970.00 |
| Dataset | Original | Guassian | Convolve | TimeGAN | ADA | ReAugment | ||||||
| MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | |
| ETTh1 | 0.405 | 0.387 | 0.407 | 0.392 | 0.416 | 0.399 | 0.409 | 0.390 | 0.407 | 0.391 | 0.3960.01 | 0.3810.01 |
| ETTh2 | 0.350 | 0.301 | 0.352 | 0.307 | 0.356 | 0.303 | 0.348 | 0.299 | 0.347 | 0.297 | 0.3460.01 | 0.2940.01 |
| ETTm1 | 0.377 | 0.341 | 0.374 | 0.340 | 0.387 | 0.352 | 0.392 | 0.357 | 0.372 | 0.336 | 0.3640.01 | 0.3280.01 |
| ETTm2 | 0.272 | 0.186 | 0.272 | 0.187 | 0.275 | 0.188 | 0.279 | 0.190 | 0.273 | 0.188 | 0.2630.01 | 0.1790.00 |
| Weather | 0.219 | 0.178 | 0.227 | 0.187 | 0.265 | 0.210 | 0.219 | 0.177 | 0.222 | 0.180 | 0.2060.00 | 0.1700.00 |
| Electricity | 0.239 | 0.148 | 0.243 | 0.150 | 0.264 | 0.170 | 0.276 | 0.183 | 0.241 | 0.149 | 0.2360.01 | 0.1470.01 |
| Traffic | 0.269 | 0.392 | 0.269 | 0.394 | 0.283 | 0.407 | 0.296 | 0.412 | 0.268 | 0.391 | 0.2640.01 | 0.3880.01 |
| Exchange | 0.206 | 0.086 | 0.208 | 0.087 | 0.210 | 0.087 | 0.210 | 0.087 | 0.209 | 0.087 | 0.2040.00 | 0.0850.00 |
| Dataset | Gaussian | Convolve | TimeGAN | ADA | ReAugment | |||||
| ETTh1 | -10.3% | -20.8% | -24.1% | -25% | -34.5% | -33.3% | -3.4% | -8.3% | 41.4% | 33.3% |
| ETTh2 | -25% | -5.3% | -16.7% | -15.8% | -33.3% | -36.8% | -50.0% | -57.9% | 191.7% | 94.7% |
| ETTm1 | 3.2% | 0.8% | 22.2% | -7.0% | 15.9% | -10.1% | 17.5% | -10.8% | 47.6% | 26.4% |
| ETTm2 | -10.0% | 0.0% | -40.0% | -16.7% | -30.0% | -11.1% | -20.0% | -16.7% | 70.0% | 44.4% |
| Weather | -75.0% | -100.0% | -183.3% | -188.9% | -66.7% | -44.4% | -125% | -122.2% | 16.7% | 22.2% |
| Elec. | -26.3% | -10.0% | -21.1% | -10.0% | -47.4% | -45.0% | -36.8% | -15.0% | 21.1% | 15.0% |
| Traffic | -2.0% | -1.4% | -4.1% | 4.1% | 6.1% | 23.0% | 0.0% | 13.5% | 51.5% | 50.0% |
| Exc. | -4.5% | -5.9% | 9.1% | 23.5% | 9.1% | 17.6% | -31.8% | -76.5% | 18.2% | 35.3% |
@@ 5.2 Results of Few-Shot Time Series Forecasting
Under the few-shot setup, we evaluate the effectiveness of ReAugment on various forecasting models, including iTransformer [21], PatchTST [23], and DLinear [38]. As shown in Table 1, our method consistently improves prediction accuracy across various public datasets when applied to different forecasting models, demonstrating its adaptability and effectiveness in automated data augmentation. Moreover, the proposed reward formulation in the REINFORCE stage further enhances performance, showing robustness across model zoo configurations. Therefore, ReAugment serves not just as a fixed augmentation scheme, but as a flexible design paradigm adaptable to different tasks and objectives. Table 2 compares the performance of various augmentation methods, all using iTransformer as the forecasting model and tripling the size of the original training set.
@@ 5.4 Model Analyses
Evaluating the improved forecasting accuracy achieved through data augmentation provides a direct measure of our method’s effectiveness. However, it does not account for the fact that the impact of data scarcity can vary significantly across different datasets. We propose a new metric that calculates the ratio of the performance promotion achieved through data augmentation in the few-shot setup, compared to the improvement obtained using the full training set. For example, when using MSE, it can be formulated as This metric considers the nature of the dataset and provides a more comprehensive evaluation of the value of augmentation. As shown in Table 4, a larger value indicates a greater performance improvement due to data augmentation.
@@
Inspired by preliminary findings, we augment the top overfit-prone samples. What would be the impact of augmenting more or fewer samples? To investigate this, we compare augmenting different proportions of overfit-prone samples with augmenting the entire few-shot dataset. For consistency, the total amount of augmented data was kept three times the size of the original training set and applied to the iTransformer model. As shown in Table 6, directly augmenting the entire few-shot dataset was less effective than data-dependent augmentation, which aligns with our preliminary findings. Furthermore, augmenting different percentages of high-variance samples resulted in varying degrees of improvement, demonstrating that our data augmentation model can adaptively enhance overfit-prone samples, leading to better performance of the forecasting model.
Computational costs.
Table 7 reports the training time for VMAE, REINFORCE, and the forecasting model in the few-shot setting. The added time from Stages A and B is acceptable given the performance gains and remains shorter than training the forecasting models. All experiments are conducted on an NVIDIA RTX 3090 GPU.
Hyperparameters.
In Suppl. C, we conduct sensitivity analyses on the size of the model zoo () and the mask rate of VMAE, and present details of other hyperparameters in our model.
| Dataset | Stage A: VMAE | Stage B: REINFORCE | Stage C: Forecasting | |
| iTransformer | PatchTST | |||
| ETTh1 | 1min | 2min | 2min | 5min |
| Elec. | 24min | 31min | 1h 33min | 2h 24min |
| Traffic | 1h 22min | 1h 53min | 4h 27min | 6h 50min |
@@ 6 Conclusions and Limitations
In this paper, we proposed ReAugment, a novel data augmentation method driven by reinforcement learning. Key technical contributions include: (i) Identifying overfit-prone data samples that could significantly benefit from augmentation by assessing their prediction diversity across a forecasting model zoo; (ii) Training a variational generative model within an RL framework to transform these overfit-prone samples into new data points, guided by a reward function derived from the performance of the model zoo, thereby enhancing both the quality and diversity of the augmented data. ReAugment significantly boosts forecasting performance while maintaining minimal computational overhead by leveraging a learnable policy to transform the overfit-prone samples.
One unresolved issue in this study is the reliance on multiple pretrained models. The proposed method introduces additional computational overhead due to the need for training VMAE, applying the REINFORCE algorithm, and performing backtesting across the model zoo. We provide a detailed comparison of the computational costs for each training stage in the experimental section.
Supplementary Material
@@ Appendix A Overall Training Pipeline
We present the overall training pipeline in Algorithm 1.
@@ Appendix B Dataset Details
ETT consists of two hourly-level datasets (ETTh) and two -minute-level datasets (ETTm). Each of them contains factors of electricity transformers, including load and oil temperature from July 2016 to July 2018.
@@
For the standard setup, we follow the data processing method of iTransformer, dividing the dataset into training, validation, and test sets, with this partitioning aligned in chronological order.
Appendix C Sensitivity Analysis
C.1 Size of Model Zoo
In the main text, we use . As shown in Table 9, we provide an analysis of the impact of varying values under the few-shot learning setup. Experiments show that using too few models affects the variance estimation, while using too many models in a few-shot setting leads to insufficient training data for each model, preventing further improvement in the final results. We observe that the model’s performance remains stable when .
| K=3 | K=4 | K=6 | K=8 | |||||
| Dataset | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE |
| ETTh1 | 0.433 | 0.412 | 0.422 | 0.403 | 0.424 | 0.404 | 0.423 | 0.403 |
| ETTh2 | 0.343 | 0.308 | 0.339 | 0.302 | 0.337 | 0.301 | 0.337 | 0.301 |
| ETTm1 | 0.429 | 0.453 | 0.410 | 0.436 | 0.413 | 0.437 | 0.411 | 0.436 |
| ETTm2 | 0.277 | 0.198 | 0.275 | 0.196 | 0.274 | 0.195 | 0.274 | 0.194 |
| Weather | 0.231 | 0.186 | 0.229 | 0.185 | 0.229 | 0.185 | 0.229 | 0.186 |
| Electricity | 0.256 | 0.168 | 0.254 | 0.165 | 0.255 | 0.166 | 0.254 | 0.164 |
| Traffic | 0.295 | 0.436 | 0.293 | 0.429 | 0.292 | 0.427 | 0.291 | 0.426 |
| Exchange | 0.225 | 0.100 | 0.224 | 0.097 | 0.224 | 0.098 | 0.224 | 0.097 |
C.2 Mask Rate of VMAE
The final model uses a mask rate of 0.3, which is kept for all experiments in the main text. As shown in Table 10, we analyze this hyperparameter in a few-shot learning setup.
| Mask rate | 0.1 | 0.3 | 0.5 | 0.7 | ||||
| MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | |
| ETTh1 | 0.428 | 0.409 | 0.422 | 0.403 | 0.425 | 0.406 | 0.435 | 0.413 |
| ETTh2 | 0.353 | 0.314 | 0.339 | 0.302 | 0.340 | 0.303 | 0.342 | 0.305 |
| ETTm1 | 0.423 | 0.448 | 0.410 | 0.436 | 0.408 | 0.435 | 0.412 | 0.439 |
| ETTm2 | 0.277 | 0.200 | 0.275 | 0.196 | 0.275 | 0.195 | 0.276 | 0.198 |
| Weather | 0.229 | 0.186 | 0.229 | 0.185 | 0.229 | 0.185 | 0.228 | 0.184 |
C.3 Additional Hyperparameters
@@
| Training Data | iTransformer | PatchTST | DLinear | |||
| MAE | MSE | MAE | MSE | MAE | MSE | |
| ETTh1 (Raw) | 0.434 | 0.411 | 0.458 | 0.446 | 0.435 | 0.408 |
| ReAugment | 0.4220.01 | 0.4030.01 | 0.4400.02 | 0.4290.02 | 0.4220.02 | 0.3880.01 |
| ETTh2 (Raw) | 0.362 | 0.320 | 0.367 | 0.321 | 0.402 | 0.356 |
| ReAugment | 0.3390.01 | 0.3020.01 | 0.3490.01 | 0.3060.01 | 0.3690.01 | 0.3340.01 |
| ETTm1 (Raw) | 0.440 | 0.470 | 0.428 | 0.457 | 0.442 | 0.471 |
| ReAugment | 0.4100.01 | 0.4360.01 | 0.4030.02 | 0.4330.02 | 0.4310.01 | 0.4620.02 |
| ETTm2 (Raw) | 0.282 | 0.204 | 0.276 | 0.199 | 0.303 | 0.219 |
| ReAugment | 0.2750.02 | 0.1960.01 | 0.2680.01 | 0.1930.01 | 0.2970.00 | 0.2160.00 |
| Weather (Raw) | 0.231 | 0.187 | 0.232 | 0.189 | 0.277 | 0.212 |
| ReAugment | 0.2290.00 | 0.1850.00 | 0.2270.00 | 0.1860.00 | 0.2760.01 | 0.2120.00 |
| Electricity (Raw) | 0.258 | 0.168 | 0.295 | 0.200 | 0.307 | 0.215 |
| ReAugment | 0.2540.01 | 0.1650.01 | 0.2750.01 | 0.1870.01 | 0.3050.01 | 0.2160.01 |
| Traffic (Raw) | 0.318 | 0.466 | 0.327 | 0.541 | 0.452 | 0.724 |
| ReAugment | 0.2930.01 | 0.4290.01 | 0.3140.02 | 0.5210.02 | 0.4060.01 | 0.6630.02 |
| Exchange (Raw) | 0.228 | 0.103 | 0.226 | 0.103 | 0.226 | 0.104 |
| ReAugment | 0.2240.00 | 0.0970.00 | 0.2230.00 | 0.0980.00 | 0.2240.00 | 0.0990.00 |
Appendix D Full Results with Multiple Training Seeds
In the main text, for all stochastic data augmentation methods, we conducted experiments using three different random seeds and reported the mean performance. It is worth noting that in all experiments, the source of randomness stems solely from the data augmentation methods, rather than the forecasting models. Accordingly, for stochastic augmentation methods, we applied different random seeds during the augmentation process, while keeping the random seed fixed for the training of the forecasting models. Here, we provide the complete results, including both the mean and standard deviation, for the following experiments:
| Dataset | Original | Gaussian | Convolve | |||
| MAE | MSE | MAE | MSE | MAE | MSE | |
| ETTh1 | 0.434 | 0.411 | 0.4370.02 | 0.4160.02 | 0.4410.03 | 0.4170.02 |
| ETTh2 | 0.362 | 0.320 | 0.3650.01 | 0.3210.01 | 0.3640.01 | 0.3230.01 |
| ETTm1 | 0.440 | 0.470 | 0.4380.03 | 0.4690.02 | 0.4260.02 | 0.4790.03 |
| ETTm2 | 0.282 | 0.204 | 0.2830.01 | 0.2040.01 | 0.2860.01 | 0.2070.01 |
| Weather | 0.231 | 0.187 | 0.2400.01 | 0.1960.00 | 0.2530.01 | 0.2040.01 |
| Electricity | 0.258 | 0.168 | 0.2630.01 | 0.1700.01 | 0.2620.01 | 0.1700.01 |
| Traffic | 0.318 | 0.466 | 0.3190.01 | 0.4670.02 | 0.3200.01 | 0.4630.01 |
| Exchange | 0.228 | 0.103 | 0.2290.00 | 0.1040.00 | 0.2260.01 | 0.0990.00 |
| Dataset | TimeGAN | ADA | ReAugment | |||
| ETTh1 | 0.4440.02 | 0.4190.01 | 0.4350.01 | 0.4130.01 | 0.4220.01 | 0.4030.01 |
| ETTh2 | 0.3660.01 | 0.3270.01 | 0.3680.00 | 0.3310.00 | 0.3390.01 | 0.3020.01 |
| ETTm1 | 0.4300.03 | 0.4830.03 | 0.4290.01 | 0.4840.01 | 0.4100.01 | 0.4360.01 |
| ETTm2 | 0.2850.02 | 0.2060.01 | 0.2840.01 | 0.2070.01 | 0.2750.02 | 0.1960.01 |
| Weather | 0.2390.01 | 0.1910.01 | 0.2460.00 | 0.1980.00 | 0.2290.00 | 0.1850.00 |
| Electricity | 0.2670.01 | 0.1770.01 | 0.2650.01 | 0.1710.00 | 0.2540.01 | 0.1650.01 |
| Traffic | 0.3150.02 | 0.4490.03 | 0.3180.01 | 0.4560.01 | 0.2930.01 | 0.4290.01 |
| Exchange | 0.2260.01 | 0.1000.01 | 0.2350.00 | 0.1160.00 | 0.2240.00 | 0.0970.00 |
| Dataset | Original | Gaussian | Convolve | |||
| MAE | MSE | MAE | MSE | MAE | MSE | |
| ETTh1 | 0.405 | 0.387 | 0.4070.02 | 0.3920.01 | 0.4160.03 | 0.3990.02 |
| ETTh2 | 0.350 | 0.301 | 0.3520.01 | 0.3070.01 | 0.3560.01 | 0.3030.01 |
| ETTm1 | 0.377 | 0.341 | 0.3740.02 | 0.3400.02 | 0.3870.02 | 0.3520.02 |
| ETTm2 | 0.272 | 0.186 | 0.2720.01 | 0.1870.00 | 0.2750.01 | 0.1880.01 |
| Weather | 0.219 | 0.178 | 0.2270.01 | 0.1870.01 | 0.2650.01 | 0.2100.01 |
| Electricity | 0.239 | 0.148 | 0.2430.01 | 0.1500.00 | 0.2640.01 | 0.1700.01 |
| Traffic | 0.269 | 0.392 | 0.2690.01 | 0.3940.02 | 0.2830.02 | 0.4070.02 |
| Exchange | 0.206 | 0.086 | 0.2080.00 | 0.0870.00 | 0.2100.01 | 0.0870.00 |
| Dataset | TimeGAN | ADA | ReAugment | |||
| ETTh1 | 0.4090.02 | 0.3900.02 | 0.4070.01 | 0.3910.01 | 0.3960.01 | 0.3810.01 |
| ETTh2 | 0.3480.01 | 0.2990.01 | 0.3470.00 | 0.2970.00 | 0.3460.01 | 0.2940.01 |
| ETTm1 | 0.3920.02 | 0.3570.03 | 0.3720.01 | 0.3360.02 | 0.3640.01 | 0.3280.01 |
| ETTm2 | 0.2790.01 | 0.1900.01 | 0.2730.00 | 0.1880.00 | 0.2630.01 | 0.1790.00 |
| Weather | 0.2190.01 | 0.1770.01 | 0.2220.00 | 0.1800.00 | 0.2060.00 | 0.1700.00 |
| Electricity | 0.2760.02 | 0.1830.01 | 0.2410.01 | 0.1490.01 | 0.2360.01 | 0.1470.01 |
| Traffic | 0.2960.02 | 0.4120.03 | 0.2680.02 | 0.3910.02 | 0.2640.01 | 0.3880.01 |
| Exchange | 0.2100.01 | 0.0870.00 | 0.2090.00 | 0.0870.00 | 0.2040.00 | 0.0850.00 |